good day my given problem is this
To produce 100 liters that is 17% acid, Liza took some amount from a 40% acid solution and a 60% acid solution then added 70 liters of pure water to the resulting mixture. How much of the 40% acid solution was used?
thank you in advance :)

Answer by rothauserc(4718)

To produce 100 liters that is 17% acid, Liza took some amount from a 40% acid solution and a 60% acid solution then added 70 liters of pure water to the resulting mixture. How much of the 40% acid solution was used?
100 * .17 = 17 liters of acid
100 - 70 = 30 liters of the 60% and 40% solutions
let x be the amount of 40% solution, then 30-x is 60%solution, so we have
.40x + .60(30-x) = 17
.40x + 18 - .60x = 17
-.20x = -1
(1/5)x = 1
x = 5
so 5 liters of the 40% solution was used